### v1

@@ 6 Discussion
In this work, we study PAC learning guarantees for precision and recall. There are two natural open questions.
First, there is a gap between the upper and lower bounds. For the scalar-loss objective, we demonstrate that an $\alpha=5$ approximate optimal scalar loss is achievable, while $\alpha=1.05$ is not, leaving it unclear what the optimal $\alpha$ is. For the Pareto-loss objective, we establish an upper bound of $(p,r)\Rightarrow(5(p+r),r)$ and a lower bound of $(p,r)\not\Rightarrow(p+0.01,r+0.01)$, again suggesting a gap that we do not yet know how to close.
Second, it remains an open question whether there exists a combinatorial measure, similar to the VC dimension in standard PAC learning, that characterizes the learnability of precision and recall. Each graph implicitly defines a distribution at each node—specifically, a uniform distribution over its neighborhood. In Section 5.3, we also link the scalar loss to the total variation distance, thus reducing the scalar loss learning problem to a special case of distribution learning. However, as shown in [lechner2024inherent], there is no such a dimension characterizing the sample complexity of learning certain distribution classes (in their case, a mixture of point mass and uniform distributions). This result suggests a potential limitation in identifying a combinatorial measure for our learning problem.

### v2

@@ 6 Discussion
One of the main goals of this paper is to study the Pareto front of precision and recall, which is reflected in our Pareto-loss objective. There are other scalar measures of predictive performance, the $F_{1}$ and the $F_{\beta}$ scores. $F_{1}$ and $F_{\beta}$ are direct functions of precision and recall, which is weaker than our Pareto-loss objective, and thus, the hardness results can be extended directly. We also note that in the realizable setting, we can minimize other alternatives for scalar losses as we can optimize both precision and recall simultaneously. Nevertheless, hypotheses with optimal $F_{1}$ or $F_{\beta}$ scores are specific points on this Pareto front, and it will be interesting for future work to find them efficiently.
Beyond that, there are two natural open questions. First, there is a gap between the upper and lower bounds. For the scalar-loss objective, we demonstrate that an $\alpha=5$ approximate optimal scalar loss is achievable, while $\alpha=1.05$ is not, leaving it unclear what the optimal $\alpha$ is. For the Pareto-loss objective, we establish an upper bound of $(p,r)\Rightarrow(5(p+r),r)$ and a lower bound of $(p,r)\not\Rightarrow(p+0.01,r+0.01)$, again suggesting a gap that we do not yet know how to close.
Second, it remains an open question whether there exists a combinatorial measure, similar to the VC dimension in standard PAC learning, that characterizes the learnability of precision and recall. Each hypothesis implicitly defines a distribution at each node—specifically, a uniform distribution over its neighborhood. In Section 5.3, we also link the scalar loss to the total variation distance, thus reducing the scalar loss learning problem to a special case of distribution learning. However, as shown in [LB24], there is no such a dimension characterizing the sample complexity of learning certain distribution classes (in their case, a mixture of point mass and uniform distributions). This result suggests a potential limitation in identifying a combinatorial measure for our learning problem.